OpenStudy (anonymous):

Find the area of the trapezoid. A. 675 ft^2 B. 337.5 ft^2 C. 187.5 ft^2 D. 150 ft^2

OpenStudy (anonymous):



OpenStudy (mertsj):

add the two bases, divide by two and multiply by the height.

OpenStudy (hoblos):

area of the trapezoid = h*(b1+b2)/2 = 15(45)/2 = 337.5 ft^2